Balance the following incomplete chemical reaction equation by inserting the proper coefficient number in front of each chemical:

__a__AgNO3(aq) + __b__CaI2(aq) → __c__Ca(NO3)2(aq) + __d__AgI(s)

Answer in the format of "a,b,c,d", e.g., "1,3,3,1".
Answer #1

We have to balance the following chemical reaction:

a AgNO3 (aq) + b CaI2 (aq) = c Ca(NO3)2 (aq) + d AgI (s)

Firstly check if the oxygen and iodine atom is balanced.

In LHS we have 2 iodine atom in CaI2 whereas in RHS we have 1 iodine atom in AgI.

So, first add coefficient 2 before AgI in RHS.

a AgNO3 (aq) + b CaI2 (aq) = c Ca(NO3)2 (aq) + 2 AgI (s)

Now, we have 2 Ag atom in RHS after adding coefficient 2 before AgI. So add coefficient 2 in LHS before AgNO3.

2 AgNO3 (aq) + b CaI2 (aq) = c Ca(NO3)2 (aq) + 2 AgI (s)

Now check the O and N atom in LHS and RHS.

In LHS we have 6 O, 2 N and in RHS we also have 6 O, 2 N.

So, final balanced reaction ill be:

2 AgNO3 (aq) + CaI2 (aq) = Ca(NO3)2 (aq) + 2 AgI (s)

So, ans is '2,1,1,2'.